The greater of two numbers is 5 more than the smaller. If the smaller is added to twice the greater, the results is 22. Find both of the numbers
Word problem always stump me
x+5 + 2x = 22
3x+5 = 22

Answer by sowmya(32)

Let the two numbers be x and y
Let x be the greater number and y be the smaller number
x = y+5
y+2x = 22
so u have to solve this system of linear equations
x-y = 5
2x + y = 22
adding,
3x = 27
x = 9
y =4
the numbers are 9 and 4